OpenStudy (anonymous):

"If f(x) = -2 - 3 and g(x) = x^2 + 5x, find each value." Problem: g(-3) If someone could answer and somehow explain how they figured it out that'd be great!

jimthompson5910 (jim_thompson5910):

g(x) = x^2 + 5x g(-3) = (-3)^2 + 5(-3) ... replace EVERY copy of x with -3 g(-3) = ???

jimthompson5910 (jim_thompson5910):

Tell me what you get

OpenStudy (anonymous):

g=8 ?

jimthompson5910 (jim_thompson5910):

incorrect

OpenStudy (anonymous):

Oh i see what I did wrong, is the answer 24?

jimthompson5910 (jim_thompson5910):

what is (-3)^2 equal to?

OpenStudy (anonymous):

9

jimthompson5910 (jim_thompson5910):

5(-3) is equal to ???

OpenStudy (anonymous):

It equals 15

jimthompson5910 (jim_thompson5910):

no, but close

OpenStudy (anonymous):

Oooh -15 then?

jimthompson5910 (jim_thompson5910):

so that means g(-3) = (-3)^2 + 5(-3) g(-3) = 9 + (-15) g(-3) = -6

OpenStudy (anonymous):

OH I always mess up the negatives, thanks a bunch @jim_thompson5910